OpenStudy (anonymous):

Given the equation y − 4 = (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (anonymous):

Add 4 both the sides and you will have your answer..

OpenStudy (anonymous):

y = 3/4x + 12

OpenStudy (anonymous):

Standard Equation of Slope-Intercept is given by: \[\huge \color{green} {y = mx + c}\] So make the given equation in this form..

OpenStudy (anonymous):

I told you to add both the sides 4... \[4 + y - 4 = x + 8 + 4\] Solve this further and tell me what you got??

OpenStudy (anonymous):

Let me what will be the Left hand side??

OpenStudy (anonymous):

3x − 4y = −40

OpenStudy (anonymous):

What??? How??

OpenStudy (anonymous):

What is left hand side tell me...

OpenStudy (anonymous):

y = 3/4x + 10

OpenStudy (anonymous):

I am asking what is left hand side of: \[4 + y - 4 = x + 8 + 4\] Tell me what is on Left of this??

OpenStudy (anonymous):

−3/4x + y = 10

OpenStudy (anonymous):

Great... Well the answer is: \[y - \cancel{4} + \cancel{4} = x + 8 + 4\] \[y = x + 12\] is your answer..

OpenStudy (anonymous):

y=x+12 is not the answer